A brass rod is 2m long at a certain temperature. What is the length for a temperature rise of 100k.(take x=0.000018-1)


1
Expert's answer
2022-02-01T15:33:51-0500

Given:

"l_0=2\:\rm m"

"\Delta T=100\:\rm K"

"\alpha=0.000018\:\rm K^{-1}"


The thermal linear expansion equation says

"l=l_0(1+\alpha\Delta T)\\
=\rm 2\: m(1+0.000018\: K^{-1}*100\: K)=2.0036\: m"